in the formula f=9/5c+32 what is the farenheiht if the celsius is -10,0,15,100?
Answer by checkley75(3666)

F=9/5C+32
F=9/5*-10+32
F=-90/5+32
F=-18+32
F=14 ANSWER FOR C=-10
-------------------------------------
F=9/5C+32
F=9/5*0+32
F=32 FOR C=0
------------------------------------
F=9/5*C+32
F=9/5*15+32
F=135/5+32
F=27+32
F=59 FOR C=15
------------------------------------------
F=9/5C=32
F=9/5*100+32
F=900/5+32
F=180+32
F=212 FOR C=100